# STAGE: Storyboard-Anchored Generation for Cinematic Multi-shot Narrative

### v1

@@ 5.1 Training details
Our implementation is based on Qwen-Image [qwenimage] where VAE is frozen. The rank of LoRA weight is set to 64. Training is performed on 8 A800 GPUs using the Adam optimizer [kingma2014adam] at a learning rate of $1\times 10^{-4}$. Our two-stage training scheme consists of an initial 100K iterations of supervised fine-tuning, followed by an additional 20K iterations of preference alignment.
@@ 5.2 Quantitative evaluation metrics
We comprehensively evaluate STAGE across five aspects [^2] Metric details are provided in the supplementary materials.: (i) Overall video quality. We assess the perceptual quality of generated videos using the Aesthetic Quality (AQ) and Image Quality (IQ) metrics from VBench [vbench]. (ii) Text video consistency. To measure the alignment between the generated video and the input text description, we employ the Overall Consistency (OC) metric from VBench [vbench]. (iii) Intra-shot coherence. We evaluate the temporal consistency within individual shots using Subject Consistency (SC) and Background Consistency (BC) from VBench [vbench]. (iv) Inter-shot coherence. We extend VBench’s intra-shot metrics for inter-shot evaluation, defining Subject Consistency-Extra (SC-E) and Background Consistency-Extra (BC-E) metrics. (v) Inter-shot transition. We introduce Transition Vector Similarity (TVS) to measure cinematic transition quality. To further evaluate the quality of multi-shot video, we employ the VLM [gemini] to rate Overall Video Quality (OVQ), Video Text Consistency (VTC), Inter-Shot Consistency (ISC), Shot Transition Smoothness (STS) on a scale of 0 to 1 score.
@@
Method Quantitative metrics evaluation LLM evaluation AQ $\uparrow$ IQ $\uparrow$ OC $\uparrow$ SC $\uparrow$ BC $\uparrow$ SC-E $\uparrow$ BC-E $\uparrow$ TVS $\uparrow$ OVQ $\uparrow$ VTC $\uparrow$ ISC $\uparrow$ STS $\uparrow$ Comparison with state-of-the-art methods CineTrans [cinetrans] 0.5652 0.6120 0.2018 0.9437 0.9504 0.6197 0.7428 0.0455 0.7972 0.3551 0.5585 0.4931 IC-LoRA [iclora] + WanX [wanx] 0.6333 0.6951 0.2140 0.9567 0.9615 0.5319 0.7438 0.2090 0.7597 0.3897 0.4901 0.4696 StoryDiffusion [storydiffusion] + WanX [wanx] 0.6941 0.7018 0.2087 0.9456 0.9640 0.5780 0.7988 0.1441 0.5343 0.2069 0.4813 0.4575 VideoGen-of-Thought [videogenofthought] 0.7210 0.6630 0.1689 0.9599 0.9554 0.6278 0.7830 0.0966 0.8106 0.1120 0.5086 0.4507 MovieAgent [movieagent] 0.5742 0.7069 0.0711 0.9664 0.9384 0.4993 0.6473 0.0079 0.4895 0.1931 0.4511 0.4182 Ours (STAGE) 0.7689 0.7305 0.2713 0.9695 0.9685 0.6917 0.8207 0.2732 0.8929 0.6069 0.6985 0.6255 Ablation study W/o MMP 0.7344 0.7220 0.2143 0.9631 0.9592 0.6088 0.7311 0.2370 0.8835 0.3642 0.5466 0.5228 W/o DES 0.7217 0.7145 0.2488 0.9542 0.9476 0.6803 0.8124 0.2680 0.7803 0.5063 0.6552 0.5167 W/o TTS 0.7476 0.7240 0.2635 0.9613 0.9633 0.6636 0.8037 0.2195 0.8733 0.5766 0.6614 0.5111
@@ 5.3 Comparison with State-of-the-art Methods
We compare our approach with state-of-the-art multi-shot video generation methods, including end-to-end approaches (i.e., CineTrans [cinetrans]) and key-frame-based approaches (i.e., IC-LoRA [iclora] + WanX [wanx], StoryDiffusion [storydiffusion] + WanX [wanx], MovieAgent [movieagent], and VideoGen-of-Thought [videogenofthought]).
Qualitative comparisons. In Fig. 4, we present visual quality comparison results with aforementioned methods. Both CineTrans [cinetrans] and IC-LoRA [iclora] + WanX [wanx] fail to maintain consistency across multiple shots (e.g., Fig.˜4 first and second row, the appearance of the train car window and the surrounding environment in the first shot are inconsistent with those in the third, leading to a disjointed narrative). StoryDiffusion [storydiffusion] + WanX [wanx] produces an overly cartoonish style and lower overall quality (e.g., Fig.˜4 third row, the entire sequence exhibits a stylized aesthetic that is unsuitable for cinematic applications). VideoGen-of-Thought [videogenofthought] suffers from poor temporal coherence in its action sequences (e.g., Fig.˜4 fourth row, the woman’s hands are on her arm in the first shot but abruptly holds a phone with hands in the second, breaking the continuity of motion). MovieAgent [movieagent] exhibits low text fidelity (e.g., Fig.˜4 fifth row, despite the prompt is sitting on a train, the generated video depicts a woman beside the train, failing to accurately reflect the user’s intent). In contrast, our STAGE workflow successfully maintains high inter-shot consistency while ensuring shot smooth transitions, resulting in a coherent and visually superior cinematic result.
@@
Method VQE TAE SCE ITE CineTrans [cinetrans] $13.2$ $18.4$ $6.4$ $16.4$ IC-LoRA [iclora] + WanX [wanx] $24.4$ $12.8$ $13.2$ $7.2$ StoryDiffusion [storydiffusion] + WanX [wanx] $1.2$ $8.0$ $5.6$ $3.6$ VideoGen-of-Thought [videogenofthought] $2.8$ $4.4$ $0.4$ $2.0$ MovieAgent [movieagent] $0.8$ $3.2$ $1.6$ $1.2$ Ours (STAGE) $\mathbf{57.6}$ $\mathbf{53.2}$ $\mathbf{72.8}$ $\mathbf{69.6}$

### v2

@@ 5.1 Training details
Our implementation is based on Qwen-Image [50] where VAE is frozen. The rank of LoRA weight is set to 64. Training is performed on 8 A800 GPUs using the Adam optimizer [16] at a learning rate of $1\times 10^{-4}$. Our two-stage training scheme consists of an initial 100K iterations of supervised fine-tuning, followed by an additional 20K iterations of preference alignment.
@@ 5.2 Quantitative evaluation metrics
We comprehensively evaluate STAGE across five aspects [^2] Metric details are provided in the supplementary materials.: (i) Overall video quality. We assess the perceptual quality of generated videos using the Aesthetic Quality (AQ) and Image Quality (IQ) metrics from VBench [14]. (ii) Text video consistency. To measure the alignment between the generated video and the input text description, we employ the Overall Consistency (OC) metric from VBench [14]. (iii) Intra-shot coherence. We evaluate the temporal consistency within individual shots using Subject Consistency (SC) and Background Consistency (BC) from VBench [14]. (iv) Inter-shot coherence. We extend VBench’s intra-shot metrics for inter-shot evaluation, defining Subject Consistency-Extra (SC-E) and Background Consistency-Extra (BC-E) metrics. (v) Inter-shot transition. We introduce Transition Vector Similarity (TVS) to measure cinematic transition quality. To further evaluate the quality of multi-shot video, we employ the VLM [42] to rate Overall Video Quality (OVQ), Video Text Consistency (VTC), Inter-Shot Consistency (ISC), Shot Transition Smoothness (STS) on a scale of 0 to 1 score.
@@
Method Quantitative metrics evaluation LLM evaluation AQ $\uparrow$ IQ $\uparrow$ OC $\uparrow$ SC $\uparrow$ BC $\uparrow$ SC-E $\uparrow$ BC-E $\uparrow$ TVS $\uparrow$ OVQ $\uparrow$ VTC $\uparrow$ ISC $\uparrow$ STS $\uparrow$ Comparison with state-of-the-art methods CineTrans [52] 0.5652 0.6120 0.2018 0.9437 0.9504 0.6197 0.7428 0.0455 0.7972 0.3551 0.5585 0.4931 IC-LoRA [13] + WanX [45] 0.6333 0.6951 0.2140 0.9567 0.9615 0.5319 0.7438 0.2090 0.7597 0.3897 0.4901 0.4696 StoryDiffusion [60] + WanX [45] 0.6941 0.7018 0.2087 0.9456 0.9640 0.5780 0.7988 0.1441 0.5343 0.2069 0.4813 0.4575 VideoGen-of-Thought [59] 0.7210 0.6630 0.1689 0.9599 0.9554 0.6278 0.7830 0.0966 0.8106 0.1120 0.5086 0.4507 MovieAgent [51] 0.5742 0.7069 0.0711 0.9664 0.9384 0.4993 0.6473 0.0079 0.4895 0.1931 0.4511 0.4182 Ours (STAGE) 0.7689 0.7305 0.2713 0.9695 0.9685 0.6917 0.8207 0.2732 0.8929 0.6069 0.6985 0.6255 Ablation study W/o MMP 0.7344 0.7220 0.2143 0.9631 0.9592 0.6088 0.7311 0.2370 0.8835 0.3642 0.5466 0.5228 W/o DES 0.7217 0.7145 0.2488 0.9542 0.9476 0.6803 0.8124 0.2680 0.7803 0.5063 0.6552 0.5167 W/o TTS 0.7476 0.7240 0.2635 0.9613 0.9633 0.6636 0.8037 0.2195 0.8733 0.5766 0.6614 0.5111
@@ 5.3 Comparison with State-of-the-art Methods
We compare our approach with state-of-the-art multi-shot video generation methods, including end-to-end approaches (i.e., CineTrans [52]) and key-frame-based approaches (i.e., IC-LoRA [13] + WanX [45], StoryDiffusion [60] + WanX [45], MovieAgent [51], and VideoGen-of-Thought [59]).
Qualitative comparisons. In Fig. 4, we present visual quality comparison results with aforementioned methods. Both CineTrans [52] and IC-LoRA [13] + WanX [45] fail to maintain consistency across multiple shots (e.g., Fig.˜4 first and second row, the appearance of the train car window and the surrounding environment in the first shot are inconsistent with those in the third, leading to a disjointed narrative). StoryDiffusion [60] + WanX [45] produces an overly cartoonish style and lower overall quality (e.g., Fig.˜4 third row, the entire sequence exhibits a stylized aesthetic that is unsuitable for cinematic applications). VideoGen-of-Thought [59] suffers from poor temporal coherence in its action sequences (e.g., Fig.˜4 fourth row, the woman’s hands are on her arm in the first shot but abruptly holds a phone with hands in the second, breaking the continuity of motion). MovieAgent [51] exhibits low text fidelity (e.g., Fig.˜4 fifth row, despite the prompt is sitting on a train, the generated video depicts a woman beside the train, failing to accurately reflect the user’s intent). In contrast, our STAGE workflow successfully maintains high inter-shot consistency while ensuring shot smooth transitions, resulting in a coherent and visually superior cinematic result.
@@
Method VQE TAE SCE ITE CineTrans [52] $13.2$ $18.4$ $6.4$ $16.4$ IC-LoRA [13] + WanX [45] $24.4$ $12.8$ $13.2$ $7.2$ StoryDiffusion [60] + WanX [45] $1.2$ $8.0$ $5.6$ $3.6$ VideoGen-of-Thought [59] $2.8$ $4.4$ $0.4$ $2.0$ MovieAgent [51] $0.8$ $3.2$ $1.6$ $1.2$ Ours (STAGE) $\mathbf{57.6}$ $\mathbf{53.2}$ $\mathbf{72.8}$ $\mathbf{69.6}$